Consider the following scenario analysis:


Rate of Return


Probability Stocks Bonds


Recession 0.20 -5% 14%


Normal 0.60 15% 8%


Boom 0.20 25% 4%



a. Calculate the expected rate of return and standard deviation for each investment.


b. Calculate coefficient of variance of each investment


c. Which investment would you prefer based on standard deviation and CV?




1
Expert's answer
2021-11-23T17:25:35-0500

Calculations for expected rate of return ,standard deviation and coefficient of variance  for each investment.



a. Calculate the expected rate of return and standard deviation for each investment.

Expected rate of return for; stock

="(0.2\times-5\%)+(0.6\times15\%)+(0.2\times25\%)=13\%"


For bonds;

"(0.2\times14\%)+(0.6\times8\%)+(0.2\times4\%)=8.4\%"


Standard deviation for; stock

"=\sqrt {variance}\times100\%"

"=\sqrt{0.0096}\times100\%"

"=9.80\%"


For bonds;

"=\sqrt {variance}\times100\%"

"=\sqrt{0.001024}\times100\%"

"=3.20\%"


c) Based on standard deviation and CV,I would prefer investment on bonds since it has the least standard deviation and CV.


Other formulas used;

Values of x represent the probabilities of the various economies.

P(x) represents the rates of return of the various states of economies.